OpenStudy (anonymous):

Solve. x^2 + 8x – 8 = 0

OpenStudy (cwrw238):

use completing the square or the quadratic formula

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

Can you check my answer after I finished it ?

OpenStudy (cwrw238):

ok

OpenStudy (anonymous):

wait a minute...

OpenStudy (anonymous):

I got \[-8 \pm \sqrt{16}\]

OpenStudy (anonymous):

How do I simplify this ?

OpenStudy (anonymous):

the answer is 4 ?

OpenStudy (cwrw238):

x^2 + 8x – 8 = 0 x^2 + 8x = 8 (x + 4)^2 - 16 = 8 (x+4^2 = 24 x + 4 = +/- sqrt24 x = sqrt24 - 4 , -sqrt24 - 4

OpenStudy (anonymous):

My answer choice -_-

OpenStudy (anonymous):

Answer is D ?

OpenStudy (cwrw238):

ok - we need to simplify sqrt24 sqrt24 = sqrt(4*6) = sqrt4 * sqrt6 = 2 sqrt6

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

Thank you :)

OpenStudy (cwrw238):

yw